If the revenue last quarter was 4.0 million EUR and the expenses were 3.7 million USD, then what would the USD/EUR exchange rate have to be for the profit margin to equal 14.0%? Report your answer rounded to exactly four decimal places. If the value is 2.104 or 2.10403, enter 2.1040 as your response. USD/EUR:1.1567
Answer #1

Profit margin = 14%

Net Income/Revenue = 0.14

Net Income = Revenue * 0.14

Net Income = 4 * 0.14 = 0.56 million EUR

Net Income = Revenue - Expenses

0.56 million = 4 - Expenses

Expenses = 4 - 0.56

Expenses = 3.44 million EUR

We need to find the exchange rate such that this expenses in EUR = 3.7 million USD

3.44 EUR = 3.7 USD

USD/EUR = 3.7/3.44

USD/EUR = 1.0755813953 ~ 1.0756 USD/EUR